### v1

@@ 2.1 Experts’ opinions and agent’s problem
${\mathbb{E}}\left[\exp\left\{\tfrac{1}{2}\int_{0}^{T}|\lambda^{(k)}(t,X_{t})|^{2}\,{\rm d}t\right\}\right]<\infty\,,$
where $\lambda^{(k)}$ is given in the definition of ${\mathcal{Q}}$.
@@
We may therefore rewrite optimisation problem (P) in terms of an optimisation over $\theta$ as follows
@@
subject to ${\mathbb{E}}^{\mathbb{Q}}[\int_{0}^{T}g_{u}\,{\rm d}u]=0$ and ${\mathbb{E}}^{\mathbb{Q}}[f_{T}]=0$, where for simplicity of notation we write $g_{u}:=g(u,X_{u})$ and $f_{T}:=f(X_{T})$, $\Delta\theta^{(k)}_{t}:=\theta(t,X_{t})-\mu^{(k)}(t,X_{t})$, and $\Sigma(t,x):=\sigma(t,x)^{\intercal}\sigma(t,x)$.
2.2 Optimal model, barycentre model, and average drift model
@@
$L_{\eta}[\theta]$ $(t,x):={\mathbb{E}}^{{\mathbb{Q}}[\theta]}_{t,x}\left[\int_{t}^{T}\tfrac{1}{2}\sum_{k\in{\mathcal{K}}}\pi_{k}\,\Delta\theta_{u}^{(k)\intercal}\,\Sigma(u,X_{u})^{-1}\Delta\theta_{u}^{(k)}\,{\rm d}u+\eta_{0}\,\int_{t}^{T}g_{u}\,{\rm d}u+\eta_{1}\,f_{T}\,\right],$
where ${\mathbb{E}}_{t,x}^{\mathbb{Q}}[\cdot]$ denotes expectation conditional on $X_{t}=x$. If we set $\eta=(0,0)$, then the problem reduces to finding the (pure) barycentre of the experts’ models, i.e., without any constraints, which we discuss in Definition 2.4 and Proposition 2.5.
@@
The following probability measure, which averages the drifts of the experts, plays an important role in the solution to optimisation problem (P).
Definition 2.1 (Average drift measure).
We define the probability measure ${\mathbb{Q}}[{\bar{\mu}}]\in{\mathcal{Q}}$, termed the “average drift measure”, which has drift ${\bar{\mu}}(t,x)$ given by
(4) ${\bar{\mu}}(t,x):=\displaystyle\sum_{k\in{\mathcal{K}}}\pi_{k}\,\mu^{(k)}(t,x)\,.$
We make the following assumptions, which is central to the existence of a solution to optimisation problem (P).
Assumption 2.3.
It holds that

### v2

@@ 2.1 Experts’ opinions and agent’s problem
$\qquad{\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}\left[\exp\left\{\tfrac{1}{2}\int_{0}^{T}|{\bar{\lambda}}_{t}|^{2}\,{\rm d}t\right\}\right]<\infty,\text{ and }{\mathbb{E}}^{{\mathbb{Q}}[\theta]}\left[\sup_{t\in[0,T]}|X_{t}|^{2}\right]<\infty\Bigg{\}}\,.$
By Theorem 1 in [22], the condition ${\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}\left[e^{\frac{1}{2}\int_{0}^{T}|{\bar{\lambda}}_{t}|^{2}\,{\rm d}t}\right]<\infty$ in the set ${\mathcal{Q}}$ ensures that ${\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}\left[\frac{{\rm d}{\mathbb{Q}}[\theta]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}\right]=1$.
Figure 1 provides a visualisation of problem (P). The left panel shows that we first change to the mean drift measure ${\mathbb{Q}}[{\bar{\mu}}]$, then to a measure ${\mathbb{Q}}[\theta]$ with arbitrary drift. The middle panel shows that we measure the KL divergence between the experts’ measures $({\mathbb{P}}^{(k)})_{k\in{\mathcal{K}}}$ and ${\mathbb{Q}}[\theta]$, and the right panel shows that we aim to have this measure be within the set of expectation constraints ${\mathcal{C}}$ in problem (P).
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{1}}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{2}}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{3}}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{4}}}$${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}{{\mathbb{Q}}[\overline{\mu}]}}$$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{\mathbb{Q}}[\theta]}$
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{1}}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{2}}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{3}}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{4}}}$$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{\mathbb{Q}}[\theta]}$
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{1}}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{2}}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{3}}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{4}}}$$\mathcal{C}$$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{\mathbb{Q}}[\theta]}$
Figure 1: A representation of problem (P), where ${\mathcal{C}}$ denotes the set of measures that attain the problem constraints.
Note that if expert-k’s model satisfies ${\mathbb{E}}^{{\mathbb{P}}^{(k)}}\big{[}\sup_{t\in[0,T]}|X_{t}|^{2}\big{]}<\infty$, then ${\mathbb{P}}^{(k)}\in{\mathcal{Q}}$. Similarly, if the average drift measure fulfils ${\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}\big{[}\sup_{t\in[0,T]}|X_{t}|^{2}\big{]}<\infty$, then ${\mathbb{Q}}[{\bar{\mu}}]\in{\mathcal{Q}}$.
2.2 Value function
Next, we rewrite the agent’s optimisation problem (P), which is an optimisation problem over probability measures, as an optimisation problem over admissible drifts of $X$. For this we define the set of admissible drifts $\theta$ of $X$ as the set of processes which induce a probability measure ${\mathbb{Q}}[\theta]\in{\mathcal{Q}}$. That is the set of admissible drifts is given by
${\mathcal{A}}:=\left\{\left.\theta:\Omega\times[0,T]\to{\mathds{R}}^{d}\;\right|\;{\mathbb{Q}}[\theta]\in{\mathcal{Q}}\right\}.$
Thus, the set ${\mathcal{A}}$ is the set of processes that generate the measures in the set ${\mathcal{Q}}$. Moreover, the set ${\mathcal{A}}$ and ${\mathcal{Q}}$ are isomorphic, one set is parametrised by ${\mathcal{F}}$-adapted processes, while the other is parametrised by probability measures. Furthermore, by Girsanov’s Theorem, $W_{t}=\int_{0}^{t}{\bar{\lambda}}_{u}\,{\rm d}u+{\overline{W}}_{t}$ is a ${\mathbb{Q}}[\theta]$-Brownian motion for each $k\in{\mathcal{K}}$, and therefore the ${\mathbb{Q}}[\theta]$-drift of $X$ is equal to $\theta$.
To rewrite optimisation problem (P) as an optimisation problem over the admissible strategies ${\mathcal{A}}$, we calculate the KL divergence from ${\mathbb{Q}}[\theta]$ to ${\mathbb{P}}^{(k)}$, and in particular the Radon-Nikodym derivative $\frac{{\rm d}{\mathbb{Q}}[\theta]}{{\rm d}{\mathbb{P}}^{(k)}}$. For any $\theta\in{\mathcal{A}}$, equivalently for any ${\mathbb{Q}}[\theta]\in{\mathcal{Q}}$, we have for all $k\in{\mathcal{K}}$, that
$\frac{{\rm d}{\mathbb{Q}}[\theta]}{{\rm d}{\mathbb{P}}^{(k)}}$ $=\frac{{\rm d}{\mathbb{Q}}[\theta]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}\;\frac{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}{{\rm d}{\mathbb{P}}^{(k)}}$
$=e^{-\frac{1}{2}\int_{0}^{T}|{\bar{\lambda}}_{t}|^{2}{\rm d}t-\int_{0}^{T}{\bar{\lambda}}^{\intercal}_{t}\,{\rm d}{\overline{W}}_{t}}\;e^{-\frac{1}{2}\int_{0}^{T}|{\bar{\gamma}}_{t}^{(k)}|^{2}{\rm d}t-\int_{0}^{T}{\bar{\gamma}}^{(k)\intercal}{\rm d}W^{(k)}_{t}}$
$=e^{-\frac{1}{2}\int_{0}^{T}(|{\bar{\lambda}}_{t}|^{2}+|{\bar{\gamma}}_{t}^{(k)}|^{2}){\rm d}t-\int_{0}^{T}{\bar{\lambda}}^{\intercal}_{t}\,({\rm d}W_{t}^{(k)}+{\bar{\gamma}}^{(k)}_{t}{\rm d}t)-\int_{0}^{T}{\bar{\gamma}}^{(k)\intercal}{\rm d}W^{(k)}_{t}}$
$=e^{-\frac{1}{2}\int_{0}^{T}(|{\bar{\lambda}}_{t}|^{2}+2{\bar{\lambda}}^{\intercal}_{t}\,{\bar{\gamma}}^{(k)}_{t}+|{\bar{\gamma}}_{t}^{(k)}|^{2}){\rm d}t-\int_{0}^{T}\big{(}\sigma_{t}^{-1}(\mu_{t}^{(k)}-\theta_{t})\big{)}^{\intercal}\,{\rm d}W^{(k)}_{t}}$
$=e^{-\frac{1}{2}\int_{0}^{T}(\mu_{t}^{(k)}-\theta_{t})^{\intercal}\Sigma_{t}^{-1}(\mu_{t}^{(k)}-\theta_{t})\,{\rm d}t-\int_{0}^{T}\big{(}\sigma_{t}^{-1}(\mu_{t}^{(k)}-\theta_{t})\big{)}^{\intercal}\,{\rm d}W^{(k)}_{t}}.$
[…]
${\mathcal{H}}\left[{\mathbb{Q}}[\theta]\,\big{\|}\,{\mathbb{P}}^{(k)}\right]$ $={\mathbb{E}}^{{\mathbb{Q}}[\theta]}\left[\log\frac{{\rm d}{\mathbb{Q}}[\theta]}{{\rm d}{\mathbb{P}}^{(k)}}\right]$
$={\mathbb{E}}^{{\mathbb{Q}}[\theta]}\left[-\tfrac{1}{2}\int_{0}^{T}(\mu_{t}^{(k)}-\theta_{t})^{\intercal}\,\Sigma_{t}^{-1}\,(\mu_{t}^{(k)}-\theta_{t}){\rm d}t-\int_{0}^{T}\big{(}\sigma_{t}^{-1}(\mu_{t}^{(k)}-\theta_{t})\big{)}^{\intercal}\,{\rm d}W_{t}^{(k)}\right]$
$={\mathbb{E}}^{{\mathbb{Q}}[\theta]}\left[-\tfrac{1}{2}\int_{0}^{T}(\mu_{t}^{(k)}-\theta_{t})^{\intercal}\,\Sigma_{t}^{-1}\,(\mu_{t}^{(k)}-\theta_{t}){\rm d}t\right.$
$\hskip 50.00008pt\left.-\int_{0}^{T}\big{(}\sigma_{t}^{-1}(\mu_{t}^{(k)}-\theta_{t})\big{)}^{\intercal}\,\left(\big{(}{\rm d}W_{t}^{(k)}+\lambda^{(k)}_{t}\,dt\big{)}-\lambda^{(k)}_{t}\,dt\right)\right]$
$={\mathbb{E}}^{{\mathbb{Q}}[\theta]}\left[\tfrac{1}{2}\int_{0}^{T}\Delta\theta^{(k)\intercal}_{t}\,\Sigma_{t}^{-1}\,\Delta\theta^{(k)}_{t}\;{\rm d}t\right],$
where, we introduce the notation $\Delta\theta^{(k)}(t,x):=\theta(t,x)-\mu^{(k)}(t,x)$, $\Delta\theta^{(k)}_{t}:=\Delta\theta^{(k)}(t,X_{t})$, $\Sigma(t,x):=\sigma(t,x)\sigma(t,x)^{\intercal}$, and $\Sigma_{t}:=\Sigma(t,X_{t})$.
@@ 2.2 Value function
Therefore, we rewrite optimisation problem (P) in terms of an optimisation over $\theta$ as follows
@@
subject to ${\mathbb{E}}^{{\mathbb{Q}}[\theta]}[\int_{0}^{T}g_{u}\,{\rm d}u]=0$ and ${\mathbb{E}}^{{\mathbb{Q}}[\theta]}[f_{T}]=0$, where for simplicity of notation we write $g_{u}:=g(u,X_{u})$ and $f_{T}:=f(X_{T})$.
@@
$L_{\eta}[\theta]$ $(t,x):={\mathbb{E}}^{{\mathbb{Q}}[\theta]}_{t,x}\left[\int_{t}^{T}\tfrac{1}{2}\sum_{k\in{\mathcal{K}}}\pi_{k}\,\Delta\theta_{u}^{(k)\intercal}\,\Sigma_{u}^{-1}\Delta\theta_{u}^{(k)}\,{\rm d}u+\eta_{0}\,\int_{t}^{T}g_{u}\,{\rm d}u+\eta_{1}\,f_{T}\,\right],$
where ${\mathbb{E}}_{t,x}^{\mathbb{Q}}[\cdot]$ denotes expectation conditional on $X_{t}=x$. If we set $\eta=(0,0)$, then the problem reduces to finding the (pure) barycentre of the experts’ models, i.e., without any constraints, which we discuss in Section 3. That is, we first discuss the pure barycentre problem and then tackle the full optimisation problem (P).
@@
Definition 2.1.
The following Mahalanobis–like distance is an important ingredient in the solution of optimisation problem (P)
(7) $\varsigma(t,x):=\tfrac{1}{2}\sum_{k\in{\mathcal{K}}}\pi_{k}\,{\Delta\mu}^{(k)}(t,x)^{\intercal}\,\Sigma^{-1}(t,x)\,{\Delta\mu}^{(k)}(t,x),$
where ${\Delta\mu}^{(k)}(t,x):=\mu^{(k)}(t,x)-{\bar{\mu}}(t,x)$.
3 The pure barycentre
In this section, we address the pure barycentre problem, i.e., the case when the agent has no additional beliefs that they wish to incorporate.
Definition 3.1 (Barycentre measure).
Consider optimisation problem (P) without any constraints, whose solution we call the barycentre (probability) measure and denote by ${\mathbb{Q}}^{\mathfrak{b}}$, that is
This problem is equivalent to solving (6) with $\eta=(0,0)$. The theorem below provides one of our key results, i.e., the drift under the barycentre measure, given in (9), and the associated value function, given in (8).
Theorem 3.2 (Barycentre Drift and Value Function).
Let Assumptions 2.1 and 2.2 be enforced. Define the function $L_{0}:[0,T]\times{\mathds{R}}^{d}\to{\mathds{R}}$, s.t.